OpenStudy (anonymous):

solve the equation please 6=-10- p _ 5 p=??????

OpenStudy (anonymous):

\[6 = -10 - \frac{p}{5}\]\[6 + 10 = -10 + 10 - \frac{p}{5}\]\[16 = -\frac{p}{5}\]\[p = 16 \times -5 = \space?\]

OpenStudy (anonymous):

THANKYOU!!!!!!! hahah

OpenStudy (anonymous):

np :)